.A researcher wants to know if Psychology majors and English majors have different opinions about poems about kittens. The researcher randomly selects 12 Psychology majors and 16 English majors. He has each person read a poem about a kitten and rate it on a scale of 0 (terrible) to 20 (fantastic). The researcher observes the following:

Mean rating for Psychology majors = 7.88 with a standard deviation of 2.37

Mean rating for English majors = 10.21 with a standard deviation of 3.15

Assume that you are working at the .05 level of significance.

  1. Write the null and alternative hypotheses (in symbols, referring to the population means).
  1. Identify the critical value(s).
  1. Calculate the t statistic (the appropriate standard error that you need in the denominator when you compute t is 1.06).
  1. State your conclusion with respect to the null hypothesis.
  1. State your conclusion with respect to the alternative hypothesis (research question).
  1. Compute Cohen’s d as an estimate of the effect size.
Answer #1

Ans:

a)

b)

df=12+16-2=26

t*=tinv(0.05,26)=+/-2.056

c)

Test statistic:

t=(7.88-10.21)/1.06

t=-2.198

d)Reject the null hypothesis,as test statistic falls in the rejection region.

e)There is sufficient evidnence to conclude that Psychology majors and English majors have different opinions about poems about kittens.

f)

Cohen's d=(7.88-10.21)/sqrt(2.37^2+3.15^2)=0.59